# eprint Nisho-2-2025

### v2

@@ Abstract
A dark matter axion with mass $m_{a}$ induces an oscillating electric field in a cylindrical sample placed under a magnetic field $B$ parallel to the cylinder axis. When the cylinder is made of a highly electrically conductive material, the induced oscillating current flows only at the surface. In contrast, if the cylinder is composed of a material with small conductivity, e.g. $\sigma=10^{-3}\text{eV}$, the electric current flows inside the bulk of the cylinder. Within the QCD axion model, the current $I$ is estimated as $I(m_{a}=10^{-4}\text{eV})\simeq 1.9\times 10^{-14}\text{A}g_{\gamma}(R/6\text{cm})^{2}(\sigma/10^{-3}\text{eV})(B/10\text{T})(10/\epsilon)(\rho_{a}/0.3\text{GeVcm}^{-3})$ with radius $R$, length $L$, permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}$ is model dependent parameter; $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$. Using an LC circuit with a quality factor $Q=10$, the signal-to-noise ratio at temperature T=100T=100mK and observation time $t_{ob}=60$second is $\sim 0.5(Q/10)\sqrt{(m_{a}/10^{-4}\text{eV})(100\text{mK}/T)(t_{ob}/60\text{s})}$. The detection of dark matter axions is feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
@@
In this letter we propose a new method for axion detection using a cylindrical sample with small electrical conductivity, $\sigma=10^{-3}\text{-}10^{-2}\text{eV}$ at low temperature 100100mK. Our target is the mass range $m_{a}=10^{-4}\mbox{-}10^{-3}$eV, which is difficult to detect with resonant cavity.
@@
On the other hand, the two suppression factors do not arise for electromagnetic fields interacting with axion, Especially, for the cylinder sample with small conductivity, the induced current mainly flows the bulk, and the resulting electric current $I$ becomes large enough to be detectable at low temperatures ( $T\sim 100\text{mK}$ ). We show that when the axion mass $m_{a}=10^{-4}$eV,
@@
$I\simeq 2.1\times 10^{-15}\text{A}g_{\gamma}\Big(\frac{R}{2\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{10\text{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (1)
with $\sigma=10m_{a}=10^{-3}$eV, where $R$ denotes the radius of the cylinder with permittivity $\epsilon=10$. The current is maximized by choosing electrical conductivity $\sigma=\epsilon m_{a}$, for instance, $\sigma=10^{-3}\text{eV}(m_{a}/10^{-4}\text{eV})$ when $\epsilon=10$. $I$ is proportional to the surface area $R^{2}$. The oscillating current flows the entire cylindrical sample, in contrast to ordinary electromagnetic induction where currents with high frequencies are confined to the surface.
In the actual measurement we compare the current $I$ with current $I_{n}$ of thermal noise. When the noise current $I_{n}$ is smaller, the real signal current $I$ can be observed more easily. Thus, it is favorable that the noise $I_{n}$ is small. We adopt cylinder sample with small conductivity, e.g. $\sigma=10^{-3}$eV, which causes large resistance $R_{c}$ of the cylinder. Because $I_{n}\propto 1/\sqrt{R_{c}}$, it is expected that the thermal noise $I_{n}$ is small for large $R_{c}$. It is an advantage of using the sample with small conductivity $\sigma=10^{-3}$eV for the axion detection. As we show later, we have
@@
$\frac{I}{I_{n}}\simeq 0.43\times 10^{-4}g_{\gamma}\sqrt{\frac{100\rm mK}{T}}\sqrt{\frac{\sigma}{10^{-3}\text{eV}}}\sqrt{\frac{10^{-4}\text{eV}}{m_{a}}}\sqrt{\frac{L}{100\mbox{cm}}}\Big(\frac{R}{2\mbox{cm}}\Big)\Big(\frac{B_{0}}{10\mbox{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
with length $L$ of the cylinder.
By adopting an appropriate conductivity, e.g. $\sigma\sim 5\times 10^{-3}\text{eV}$ at low temperature $\sim 20$mK and radius ( length ) of the cylinder 66cm ( 100100cm ), we show that using parallel LC circuit tuned with a quality factor $Q=10$, dark matter axion can be probed in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$ with signal-noise ratio of the electric current bigger than $10$.
Our method using cylinder with small electrical conductivity is also effective for so called, dark photon. The electromagnetic coupling with the dark photon gives rise to similar effect to the one in the QCD axion that oscillating electric current is induced in the whole of cylinder.
@@
where $\vec{E}$ and $\vec{B}$ denote the electric and magnetic fields, respectively. The coupling constant is $g_{a\gamma\gamma}=g_{\gamma}\alpha/\pi f_{a}$, with the fine structure constant $\alpha\simeq 1/137$, the axion decay constant $f_{a}$, and the relation $m_{a}f_{a}\simeq 6\times 10^{-6}\text{eV}\times 10^{12}\text{GeV}$ for the QCD axion model. The model dependent coefficient is $g_{\gamma}\simeq 0.37$ for the DFSZ model dfsz ; dfsz1 and $g_{\gamma}\simeq-0.96$ for the KSVZ model ksvz ; ksvz1 .
@@
In this letter we show that, for a cylindrical sample in Fig.1 with small electrical conductivity, the electric current induced by axion dark matter can be large enough to be detectable at low temperatures $T\sim 100\text{mK}$. In this case, the current flows throughout the bulk of the cylinder, while for a good conductor with large conductivity, the current is confined to a thin surface layer.
We focus on the axion mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$, using a cylinder of length $L=100\text{cm}$, radius $R\geq 2$ cm, and conductivity $\sigma=10m_{a}=10^{-3}\text{-}10^{-2}\text{eV}$. As we show soon later, the condition of $\sigma/m_{a}=\epsilon$ maximize the electric field induced in the cylinder with permittivity $\epsilon$. Such a matter with small conductivity may be realized at low temperatures $\sim 100\text{mK}$, for instance by using semiconductors with impurity doping appropriately.
@@
We remind that $y\equiv\sigma/m_{a}$ and $x\equiv m_{a}R$ and $b=b(y)$ is the function of $y$. $P$ is a complicated function in $x$ and $y$, or $m_{a}$ and $R$. But we show below that $P$ becomes a simple function in $x$ and $y$ when we take the limit of large conductivity, $\sigma\sim 10^{4}$eV or small $\sigma\sim 10^{-3}$eV. We note that electric field is written such that $E^{\prime}(\rho)=E_{a}U$.
It is easy to confirm that in the limit of $\sigma\to\infty$ ( $y\to\infty$ ), $P\propto m_{a}\rho_{a}V(B_{0}/m_{a}M)^{2}|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}$ with $M\equiv\pi f_{a}/(g_{\gamma}\alpha)$ and the volume $V=2\pi RL\delta$, within which the electric current flows. $\delta=\sqrt{2/m_{a}\sigma}$ denotes skin depth. The current flows only in the surface with the depth $\delta$ of the cylinder. It is coincide with the previous result kishimoto . In the limit, the main contribution comes from the second term in $U$. We note that $|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}\simeq 1$ for $x>5$.
The skin depth for general $\sigma$ represented by the second term of $U(x,y,z)$ in eq(18) is given by $\big(m_{a}\mbox{Im}(b(y))\big)^{-1}=\Big(m_{a}(\epsilon^{2}+(\sigma/m_{a})^{2})^{1/4}\sin(\theta/2)\Big)^{-1}>\delta=\sqrt{2/m_{a}\sigma}$. In this paper, assuming typical value $\epsilon=10$ in semiconductor, we take small electric conductivity $\sigma\sim\epsilon m_{a}\sim 10m_{a}$ ( $y\sim 10$ ) and consider the mass region $m_{a}=10^{-3}\mbox{-}10^{-4}\mbox{eV}$. Then, the skin depth $\big(m_{a}\mbox{Im}(b(y))\big)^{-1}\simeq 0.1\text{cm}(10^{-4}\text{eV}/m_{a})$, smaller than the radius R=2R=2cm examined in our paper. On the other hand, the first term of $U$ is present in the whole of the cylinder, not confined at the surface.
@@
Numerically, noting $|E_{a}|=|g_{a\gamma\gamma}a(t)B_{0}|\simeq 1.6\times 10^{-36}\text{GeV}^{2}g_{\gamma}(B_{0}/10\text{T})$, we find
$P(\sigma=10^{-3}\text{eV})\simeq 2.8\times 10^{-28}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (22)
with $\epsilon=10$, $y=10$ ( $\sigma=10m_{a}$ ) and $x\simeq 10$ ( R=2R=2cm ). Obviously, $P$ in eq.(21) is proportional to $R^{2}=x^{2}/m_{a}^{2}$, that is, the cross section $\pi R^{2}$ of the cylinder. The energy dissipation takes place in the bulk volume $\propto R^{2}L$, not merely in the surface, resulting in substantially enhanced dissipation power $P$.
@@
$\simeq$ $8.0\times 10^{-32}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)\sqrt{\frac{m_{a}}{10^{-4}\mbox{eV}}}\sqrt{\frac{10^{4}\mbox{eV}}{\sigma}}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (24)
with $y=\sigma/m_{a}=10^{8}$ and $x=m_{a}R\simeq 10$. It is generated in the surface with the skin depth $\delta\sim 10^{-5}$cm of the cylinder. We find that the power $P(\sigma=10^{-3}\text{eV})$ with small conductivity, is three order of magnitude larger than $P(\sigma=10^{4}\text{eV})$ with large conductivity. That is, $P(\sigma=10^{-3}\text{eV})>10^{3}P(\sigma=10^{4}\text{eV})$. It is the reason why we use such a cylinder with small conductivity $\sigma\sim 10^{-3}$eV.
@@
$I(\sigma=10^{-3}\text{eV})=\frac{\sigma|E_{a}|S}{2\epsilon}\simeq 2.1\times 10^{-15}\text{A}g_{\gamma}\Big(\frac{R}{2\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{10\text{T}}\Big),$ (26)
@@
where the electric field $E=Re(E^{\prime})=|E_{a}|/2\epsilon=|E_{a}|/20$ is the one in the first term with $y=10$ in eq(14). We detect the axion by measuring the large electric current $I$.
@@
The point causing the large current $I$ is that electric field $E$ inside of the cylinder with small conductivity, e.g. $\sigma=10^{-3}$eV has no suppression factor $\sqrt{m_{a}/\sigma}\simeq 10^{-4}$, which is present when the cylinder with large conductivity $\sigma=10^{4}$eV. Additionally, the current flows in the whole of the cylinder, not just in the surface. The facts make the large electric current $I$ to be observable.
@@
There is another point which makes axion detection feasible. In the actual observation, we need to take into account thermal noise ( Johnson-Nyquist noise ). The electric current $I_{n}$ by the thermal noise is given by $I_{n}=\sqrt{2T\delta\omega/\pi R_{c}}$ with temperature T=100T=100mK and the width $\delta\omega$ of the frequency; $\delta\omega=10^{-6}m_{a}=10^{-10}(m_{a}/10^{-4}\text{eV})$eV. $R_{c}=L/(\sigma S)$ with $S=\pi R^{2}$ denotes electric resistance of the cylinder; $R_{c}=0.16(L/100\text{cm})(2\text{cm}/R)^{2}$. Then we find
$I_{n}=4.9\times 10^{-11}\text{A}\sqrt{\Big(\frac{T}{100\rm mK}\Big)\Big(\frac{m_{a}}{10^{-4}\rm eV}\Big)\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{100\text{cm}}{L}\Big)\Big(\frac{R}{2\text{cm}}\Big)^{2}}.$ (27)
Because the noise $I_{n}\propto 1/\sqrt{R_{c}}$, the noise is small for the large resistance $R_{c}\simeq 0.16$ of the cylinder. It should be stressed that the small noise $I_{n}$ is caused by small electrical conductivity $\sigma=10^{-3}$eV of the cylinder. Thus, we obtain the ratio of the current $I$ to the noise $I_{n}$,
$\frac{I(\sigma=10^{-3}\text{eV})}{I_{n}(\sigma=10^{-3}\text{eV})}\simeq 0.43\times 10^{-4}g_{\gamma}\sqrt{\frac{100\rm mK}{T}}\sqrt{\frac{\sigma}{10^{-3}\text{eV}}}\sqrt{\frac{10^{-4}\text{eV}}{m_{a}}}\sqrt{\frac{L}{100\mbox{cm}}}\Big(\frac{R}{2\mbox{cm}}\Big)\Big(\frac{B_{0}}{10\mbox{T}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (28)
In order to amplify the above ratio $I/I_{n}$, we use, for instance, parallel LC circuit with resistance $R_{c}$, inductance $L_{c}$ and capacitance $C_{c}$. ( The resistance $R_{c}$ in the parallel LC circuit is almost equal to the one of the cylinder. ) When we tune the circuit with resonant frequency $\omega=\sqrt{1/L_{c}C_{c}}=m_{a}$ and a quality factor $Q$ such as $Q=R_{c}\sqrt{C_{c}/L_{c}}$, the above ratio is amplified such as $Q(I/I_{n})$. In actual experiment, it is difficult to have large quality factor $Q$ such as $Q>10^{3}$. So we assume $Q=10$ as a realistic value and measure the electric current in the coil of the LC circuit. Furthermore, in the measurement of the oscillating current, we take the frequency width $\delta\omega=10^{-6}m_{a}$ and observation time $60$ seconds. Then, the actual noise is reduced as $I_{n}/\sqrt{\delta\omega\delta t_{ob}/(2\pi)}\simeq I_{n}/(1.2\times 10^{3}\sqrt{\frac{m_{a}}{10^{-4}\text{eV}}}\sqrt{\frac{t_{ob}}{60s}})$.
It follows that the signal-noise ( SN ) ratio is
@@
$\frac{I(\sigma=10^{-3}\text{eV})}{I_{n}(\sigma=10^{-3}\text{eV})}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 0.51\,g_{\gamma}\Big(\frac{Q}{10}\Big)\sqrt{\frac{100\rm mK}{T}}\sqrt{\frac{t_{ob}}{60\rm\,s}}\sqrt{\frac{\sigma}{10^{-3}\text{eV}}}\sqrt{\frac{L}{100\mbox{cm}}}\Big(\frac{R}{2\mbox{cm}}\Big)\Big(\frac{B_{0}}{10\mbox{T}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (29)
with $y=10$ and $x\simeq 10$, where $t_{ob}=60$ s denotes the observation time. Although the SN ratio does not explicitly involve the axion mass $m_{a}$, it implicitly depends on $m_{a}$ because $\sigma$ is taken as $\sigma=10m_{a}$. When we searches $m_{a}=10^{-3}$eV, it is favorable to have $\sigma=10^{-2}$eV. Then, the SN ratio is $\simeq 1.6$.
Therefore, it turns out that by the observation of the electric current $I$ with appropriate choices of, for instance, temperature T=20T=20mK and the radius R=6R=6cm, and observation time $t_{ob}=600$ s, SN ratio is larger than $10$. Without such a choice of the parameters, we can obtain large SN ratio as far as we tune quality factor $Q$ bigger than $10$.
We point out that when we search much larger axion mass $m_{a}>10^{-3}$eV, it is easier to detect such axions This is because electric current $I\propto\sigma$ and SN ratio $\propto\sqrt{\sigma}$ are much larger than those of $m_{a}=10^{-3}$eV with the appropriate choice of conductivity $\sigma=10m_{a}$.
We find that the method of the axion detection using cylinder sample with small electrical conductivity is very powerful. This is because the axion–electromagnetic coupling induces an oscillating electric field $|E_{a}|/2\epsilon$ throughout the entire cylindrical sample, producing sufficiently large electric current $I\propto\sigma(|E_{a}|/\epsilon)^{2}R^{2}$ for the detection. Furthermore, the sample with small electrical conductivity has small thermal noise compared with the one with large electrical conductivity.

### v3

@@ Abstract
A dark matter axion with mass $m_{a}$ induces an oscillating electric field in a cylindrical sample placed under a magnetic field $B_{0}$ parallel to the cylinder axis. When the cylinder is made of a highly electrically conductive material, the induced oscillating current flows only at the surface. In contrast, if the cylinder is composed of a material with small conductivity, e.g. $\sigma=10^{-3}\text{eV}$, the electric current flows inside the bulk of the cylinder. Within the QCD axion model, the power $P$ is estimated as $P\simeq 5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ with radius $R$, length $L$, permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}$ is model dependent parameter; $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$. The detection of dark matter axions in our method may be feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
I introduction
@@
In this letter we propose a new method for axion detection using a cylindrical sample with small electrical conductivity, $\sigma=10^{-3}\text{-}10^{-2}\text{eV}$ at low temperature 2020mK. Our target is the mass range $m_{a}=10^{-4}\mbox{-}10^{-3}$eV, which is difficult to detect with resonant cavity.
@@
On the other hand, the two suppression factors do not arise in materials with small conductivity such as semiconductor, Especially, for the cylinder sample with small conductivity $\sigma=10^{-3}$eV, the induced current mainly flows the bulk, and the resulting electric current $I$ and power $P$ become large. It would be detectable at low temperatures ( $T\sim 10\text{mK}$ ). We show that when the axion mass $m_{a}=10^{-4}$eV,
@@
$I$ $\simeq$ $2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (1)
@@
$P(\sigma=10^{-3}\text{eV})$ $\simeq$ $5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
with the choice $\sigma=10m_{a}=10^{-3}$eV, where $R(L)$ denotes the radius ( length ) of the cylinder with permittivity $\epsilon=10$. The power is maximized by choosing electrical conductivity such as $\sigma=\epsilon m_{a}$, for instance, $\sigma=10^{-3}\text{eV}(m_{a}/10^{-4}\text{eV})$ when $\epsilon=10$. We note that $I$ and $P(\sigma=10^{-3}\text{eV})$ are proportional to the surface area $R^{2}$.
The oscillating current flows the entire cylindrical sample, in contrast to ordinal metal with large conductivity, in which electric currents with high frequencies are confined to the surface due to electromagnetic induction.
Indeed, when the conductivity is large, for instance, $\sigma=10^{4}$eV, the power is given by
$P(\sigma=10^{4}\text{eV})\simeq 5.4\times 10^{-31}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)\sqrt{\frac{m_{a}}{10^{-4}\mbox{eV}}}\sqrt{\frac{10^{4}\mbox{eV}}{\sigma}}\Big(\frac{B}{15\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (3)
Obviously, the power $P(\sigma=10^{4}\text{eV})$ proportional to $R$ is fourth order of magnitude smaller than $P(\sigma=10^{-3}\text{eV})$.
In the observation, the power $P(\sigma=10^{-3}\text{eV})$ should be compared with thermal noise $P_{n}=T\delta\omega/2\pi$ with $\delta\omega=10^{-6}m_{a}$,
$\frac{P(\sigma=10^{-3}\text{eV})}{P_{n}}\simeq 8.6\times 10^{-7}g_{\gamma}^{2}\Big(\frac{20\text{mK}}{T}\Big)\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{B}{15\text{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (4)
The noise is further reduced by considering a frequency width $\delta\omega=10^{-6}m_{a}$ and an observation time $\delta t_{ob}=100$seconds. In this case, the noise decreases by the factor $\sqrt{\delta\omega\delta t_{ob}/2\pi}\simeq 1.6\times 10^{3}$, yielding $P(\sigma=10^{-3}\text{eV})\sqrt{\delta\omega\delta t_{ob}/2\pi}/P_{n}\simeq 1.3\times 10^{-3}$.
Although the details of power detection are beyond the scope of this paper, we expect the ratio to be sufficiently large for the power ( $P(\sigma=10^{-3},\text{eV})$ ) to become detectable in the near future. Our method, which employs a cylinder with low electrical conductivity, is also applicable to the detection of the so-called dark photon. The electromagnetic coupling of the dark photon produces an effect similar to that of the axion.
II axion
@@
where $\vec{E}$ and $\vec{B}$ denote the electric and magnetic fields, respectively. The coupling constant is $g_{a\gamma\gamma}=g_{\gamma}\alpha/\pi f_{a}$, with the fine structure constant $\alpha\simeq 1/137$, the axion decay constant $f_{a}$, and the relation $m_{a}f_{a}\simeq 6\times 10^{-3}\text{GeV}^{2}$ for the QCD axion model. The model dependent coefficient is $g_{\gamma}\simeq 0.37$ for the DFSZ model dfsz ; dfsz1 and $g_{\gamma}\simeq-0.96$ for the KSVZ model ksvz ; ksvz1 .
@@
III electric power generated in cylinder with small conductivity
In this letter we show that, for a cylindrical sample in Fig.1 with small electrical conductivity, the electric power induced by axion dark matter can be large and may be detectable at low temperatures $T\sim 10\text{mK}$. The current flows throughout the bulk of the cylinder, while for a good conductor with large conductivity, the current is confined to a thin surface layer.
We focus on the axion mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$, using a cylinder of length $L=100\text{cm}$, radius $R=6$ cm, and conductivity $\sigma=10m_{a}=10^{-3}\text{-}10^{-2}\text{eV}$. As we show soon later, the condition of $\sigma=\epsilon m_{a}$ maximize the electric power induced in the cylinder with permittivity $\epsilon$. Such a matter with small conductivity must be realized at low temperatures $\sim 10\text{mK}$, for instance by using semiconductors with impurity doping appropriately.
@@
We remind that $y\equiv\sigma/m_{a}$ and $x\equiv m_{a}R$ and $b=b(y)$ is the function of $y$. $P$ is a complicated function in $x$ and $y$, or $m_{a}$ and $R$. But we show below that $P$ becomes a simple function in $x$ and $y$ when we take the limit of large conductivity, $\sigma\sim 10^{4}$eV or small one $\sigma\sim 10^{-3}$eV. We note that electric field is written such that $E^{\prime}(\rho)=E_{a}U$.
It is easy to confirm that in the limit of $\sigma\to\infty$ ( $y\to\infty$ ), $P\propto m_{a}\rho_{a}V(B_{0}/m_{a}M)^{2}|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}$ with $M\equiv\pi f_{a}/(g_{\gamma}\alpha)=1/g_{a\gamma\gamma}$ and the volume $V=2\pi RL\delta$, within which the electric current flows. $\delta=\sqrt{2/m_{a}\sigma}$ denotes skin depth. The current flows only in the surface with the depth $\delta$ of the cylinder. It is coincident with the previous result kishimoto . In the limit, the main contribution comes from the second term in $U$. We note that $|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}\simeq 1$ for $x>5$.
The skin depth for general $\sigma$ represented by the second term of $U(x,y,z)$ in eq(20) is given by $\big(m_{a}\mbox{Im}(b(y))\big)^{-1}=\Big(m_{a}(\epsilon^{2}+(\sigma/m_{a})^{2})^{1/4}\sin(\theta/2)\Big)^{-1}>\delta=\sqrt{2/m_{a}\sigma}$. In this paper, assuming typical value $\epsilon=10$ in semiconductor, we take small electric conductivity $\sigma\sim\epsilon m_{a}\sim 10m_{a}$ ( $y\sim 10$ ) and consider the mass region $m_{a}=10^{-3}\mbox{-}10^{-4}\mbox{eV}$. Then, the skin depth $\big(m_{a}\mbox{Im}(b(y))\big)^{-1}\simeq 0.1\text{cm}(10^{-4}\text{eV}/m_{a})$, smaller than the radius R=6R=6cm examined in our paper. On the other hand, the first term of $U$ is present in the whole of the cylinder, not confined at the surface.
@@
Numerically, noting $|E_{a}|=|g_{a\gamma\gamma}a(t)B_{0}|\simeq 2.4\times 10^{-36}\text{GeV}^{2}g_{\gamma}(B_{0}/15\text{T})$, we find
$P(\sigma=10^{-3}\text{eV})\simeq 5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (24)
with $\epsilon=10$, $y=10$ ( $\sigma=10m_{a}$ ) and $x\simeq 30$ ( R=6R=6cm ). Obviously, $P(\sigma=10^{-3}\text{eV})$ in eq.(24) is proportional to $R^{2}=x^{2}/m_{a}^{2}$, that is, the cross section $\pi R^{2}$ of the cylinder. The energy dissipation takes place in the bulk volume $\propto R^{2}L$, not merely in the surface, resulting in substantially enhanced dissipation power $P$.
@@
$\simeq$ $5.4\times 10^{-31}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)\sqrt{\frac{m_{a}}{10^{-4}\mbox{eV}}}\sqrt{\frac{10^{4}\mbox{eV}}{\sigma}}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$
with $y=\sigma/m_{a}=10^{8}$ and $x=m_{a}R\simeq 30$. It is generated in the surface with the skin depth $\delta\sim 10^{-5}$cm of the cylinder. We find that the power $P(\sigma=10^{-3}\text{eV})$ is fourth order of magnitude larger than $P(\sigma=10^{4}\text{eV})$. That is, $P(\sigma=10^{-3}\text{eV})>10^{4}P(\sigma=10^{4}\text{eV})$. It is the reason why we use such a cylinder with small conductivity $\sigma\sim 10^{-3}$eV.
@@
$I(\sigma=10^{-3}\text{eV})=\frac{\sigma|E_{a}|S}{2\epsilon}\simeq 2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big),$ (27)
@@
where the electric field $E=Re(E^{\prime})=|E_{a}|/2\epsilon=|E_{a}|/20$ is the one in the first term with $y=10$ in eq(16).
@@
The point causing the large current $I(\sigma=10^{-3}\text{eV})$ is that electric field $E$ inside of the cylinder with small conductivity, e.g. $\sigma=10^{-3}$eV has no suppression factor $\sqrt{m_{a}/\sigma}\simeq 10^{-4}$, which is present when the cylinder with large conductivity $\sigma=10^{4}$eV. Additionally, the current flows in the whole of the cylinder, not just in the surface. The facts make the large electric current $I(\sigma=10^{-3}\text{eV})\sim 10^{-14}$A.
@@
In order to observe the power $P(\sigma=10^{-3}\text{eV})$, we need to take into account thermal noise ( Johnson-Nyquist noise ). The noise power $P_{n}$ is given such that $P_{n}=T\delta\omega/2\pi$ with $\delta\omega=10^{-6}m_{a}$. $\delta\omega$ denotes frequency width in the observation.
Then we have
$P_{n}=\frac{T\delta\omega}{2\pi}\simeq 6.6\times 10^{-21}\text{W}\Big(\frac{T}{20\text{mK}}\Big)\Big(\frac{m_{a}}{10^{-4}\text{eV}}\Big).$ (28)
The noise is reduced by the factor $\sqrt{\delta\omega\delta t_{ob}/2\pi}$ with observation time $\delta t_{ob}$. It yields
$P_{n}\to\frac{P_{n}}{\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}}\simeq 4.2\times 10^{-24}\text{W}\Big(\frac{T}{20\text{mK}}\Big)\Big(\frac{m_{a}}{10^{-4}\text{eV}}\Big)^{1/2}\Big(\frac{100\,\text{s}}{\delta t_{ob}}\Big)^{-1/2}.$ (29)
Therefore, the ratio of $P(\sigma=10^{-3}\text{eV})$ to the noise $P_{n}/\sqrt{\delta\omega\delta t_{ob}/2\pi}$ is
@@
$\frac{P(\sigma=10^{-3}\text{eV})}{P_{n}}\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}\simeq 1.4\times 10^{-3}g_{\gamma}^{2}\Big(\frac{20\text{mK}}{T}\Big)\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\times$
$\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)^{1/2}\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}.$ (30)
It is the ratio when $m_{a}=10^{-4}$eV and $\sigma=10^{-3}$eV. Further, when $m_{a}=10^{-3}$eV, taking $\sigma=10m_{a}=10^{-2}$eV, we have
$\frac{P(\sigma=10^{-2}\text{eV})}{P_{n}}\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}\simeq 4.4\times 10^{-3}g_{\gamma}^{2}\Big(\frac{20\text{mK}}{T}\Big)\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\times$
$\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-3}\mbox{eV}}\Big)^{1/2}\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}.$ (31)
To achieve a signal-to-noise ratio greater than 1, the power $P$ must be amplified by approximately three orders of magnitude. The method of amplification and its technical implementation are beyond the scope of this paper. However, once such amplification is realized, the dark matter axion can be detected using our proposed approach.
IV conclusion